# Tin-Based Chalcogenide Perovskites: A Promising Lead-Free Alternative for Stable and High-Performance Photovoltaics

### v1

@@ III.6 Spectroscopic Limited Maximum Efficiency:
The above discussed properties indicate that the investigated systems hold great potential for photovoltaic applications and to validate this, we also calculated their power conversion efficiency (PCE). The theoretical PCE of each system is calculated using the spectroscopic limited maximum efficiency (SLME) method. SLME was introduced by Yu and Zunger[44] (for details, see the SM), which is an improved version of the Shockley-Queisser (SQ) limit[77]. The latter is less realistic since it disregards the losses resulting from radiative recombinations due to the non-conservation of the absorbed photon momentum. SLME incorporates the magnitude of the bandgap and its nature (direct or indirect), shape of absorption spectra, thickness of the absorber layer, the material-dependent non-radiative recombination losses, and the temperature. The standard solar spectrum (AM-1.5G), the absorption coefficient, thickness, and the electronic G0W0@PBE bandgap are thus used as inputs to evaluate the theoretical SLME of ASnX3 (A = Ca, Sr, Ba; X = S, Se) CPs at 293.15 K temperature.
In addition, the optical transition possibility from VBM to CBM for these CPs has been confirmed through the computation of transition dipole moment matrix (P) elements; its square (P2) gives the transition probability between the initial (VBM) and the final (CBM) state. Figure 4(a) shows the optically allowed dipole transition at -point for CaSnS3, and for other configurations, see Figure S5. Despite having an indirect electronic bandgap, BaSnS3 and BaSnSe3 exhibit an optically allowed dipole transition at their direct band edge. However, when is not the minimum bandgap of the materials (i.e., ), non-radiative recombination plays a vital role in the SLME calculation and the radiative recombinations vary with a factor , where , is the Boltzmann constant, and T is the temperature[44] (for details, see the SM).
Next, the thickness dependence of SLME has been computed for all the ASnX3 (A = Ca, Sr, Ba; X = S, Se) compounds using BSE@G0W0@PBE method and plotted in Figure 4(b). It is clear that the SLME rises as thickness increases and eventually saturates beyond a certain thickness. The maximum SLME is found to be 31.20% for CaSnS3, which is consistent with the previously documented theoretical efficiency of 32.45% at 10 m thickness[14] and is higher than CH3NH3PbI3 (28.97% at 2 m)[28]. The highest SLME values for SrSnS3, BaSnS3, CaSnSe3, SrSnSe3, and BaSnSe3 compounds are calculated as 29.53%, 26.66%, 24.20%, 25.95%, and 26.09%, respectively. These values of SLME are favorable for the photovoltaic applications compared to ABS3 (A = Ca, Sr, Ba; B = Zr, Hf), CsGeX3 (X = Cl, Br, I), and other halide perovskites[14, 78, 79, 80, 22] (for details, see Table S13 of the SM).
@@ III.7 Analysis of SCAPS-1D results:
To verify the photovoltaic potential of the investigated systems and our calculated SLME, we also conducted a detailed analysis using the Solar Cell Capacitance Simulator in 1 Dimension (SCAPS-1D). This simulation allowed us to evaluate key performance metrics and gain deeper insights into the efficiency and operational behavior of the devices. Here, the performance of PSC is evaluated by varying the thickness of the absorber layer while keeping the ETL, HTL, and FTO thicknesses constant. The absorber layer thickness is varied from 0.01 m to 1.2 m and found that the thickness of the perovskite layer has a significant impact on the performance of the PSC [see Figure 5(b)]. After reaching the optimum efficiency, the effect of the thickness of the perovskite layer seems to saturate, in agreement to our SLME calculations. The obtained results are summarized in the Table 7. Our findings indicate that the device simulation results are consistent with the SLME work, specifically for ASnS3 (A = Ca, Sr, Ba). On the other hand, the minor discrepancies between the device simulation results for ASnSe3 (A = Ca, Sr, Ba) and the SLME findings may be attributed to the materials used in the ETL, HTL, and FTO layers. In general, SCAPS-1D proves to be a highly reliable software, as its simulation results confirm the high efficiency of the investigated materials, closely aligning with theoretical predictions.
@@
Overall, study of all the above discussed properties collectively show that Se-based CaSnSe3 and SrSnSe3 demonstrate lower exciton binding energies and higher charge carrier mobilities than the rest of the investigated systems. However, their SLME is also relatively low. On the other hand, S-based CaSnS3 and SrSnS3 perovskites exhibit higher SLME despite having higher exciton binding energies and lower charge carrier mobilities compared to the other examined materials. This improved SLME is primarily attributed to the optimal bandgaps of CaSnS3 and SrSnS3 compared to their selenium (Se) counterparts, specifically from the perspective of solar cells. This reveals a trade-off between exciton binding energy and charge carrier mobility, and the bandgap of these materials. This suggests that integrating a combination of sulfur (S) and selenium (Se) atoms in these materials could be advantageous for potential application in solar cells, meriting further investigation.
@@ IV Conclusions:
In conclusion, we have carried out a comprehensive study to investigate the ground- and excited-state properties of distorted chalcogenide perovskites (ASnX3; A = Ca, Sr, Ba, and X = S, Se) under the framework of state-of-the-art DFT combined with DFPT and MBPT (viz., GW and BSE). The mechanical properties confirm the stability, and Pugh’s and Poisson’s ratios reveal the ductile nature of these perovskites. The bandgaps calculated using G0W0@PBE method are in the range of 0.791.50 eV, and the small effective masses of electrons suggest good charge carrier mobility, which is advantageous for energy-harvesting properties. Furthermore, they have a high optical absorption coefficient ( 104 cm-1) and an optical bandgap extending from the near-infrared to the visible range, which are indicated by the BSE calculations. The exciton binding energies of these compounds, ranging from 0.04 to 0.23 eV, are similar to those of Zr- and Hf-based CPs but are little higher than those typically found in conventional lead-based HPs. Also, these perovskites have smaller carrier-phonon coupling strengths than the conventional lead-based HPs as well as Zr- and Hf-based CPs, which leads to better polaron mobility for electrons (21.33416.02 cm2V-1s-1) and holes (7.02260.69 cm2V-1s-1). Additionally, the Fröhlich mesoscopic model implies that the charge-separated polaronic states have lower stability in comparison to bound excitons. Lastly, the SLME method forecasts that one could achieve the highest PCE of up to 31.20% by employing them, which is corroborated by conventional device (FTO/TiO2/ASnX3/Spiro-OMeTAD/Au) simulations using SCAPS-1D software. Overall, these results are anticipated to expedite the research and use of chalcogenide perovskites in optoelectronic applications, in general, and solar cell technology, in particular.
@@ Acknowledgements.
The authors would like to acknowledge the Council of Scientific and Industrial Research (CSIR), Government of India [Grant No. 3WS(007)/2023-24/EMR-II/ASPIRE] for financial support. S.A. would also like to acknowledge the CSIR, India [Grant No. 09/1128(11453)/2021-EMR-I] for the Senior Research Fellowship. S. D. thanks Prof. Marc Burgelman, Department of Electronics and Information Systems (ELIS), University of Gent, for providing the SCAPS-1D software. The high performance computing facility "Magus" and workstations available at the School of Natural Sciences, Shiv Nadar Institution of Eminence (SNIoE), were used to perform all calculations.References

### v2

@@ III.6 Spectroscopic Limited Maximum Efficiency:
The above-discussed properties indicate that the investigated systems hold great potential for photovoltaic applications, and to validate this, we also calculated their power conversion efficiency (PCE). The theoretical PCE of each system is calculated using the spectroscopic limited maximum efficiency (SLME) method. SLME was introduced by Yu and Zunger[44] (for details, see the Supplemental Material), which is an improved version of the Shockley-Queisser (SQ) limit[77]. The latter is less realistic since it disregards the losses resulting from radiative recombinations due to the non-conservation of the absorbed photon momentum. SLME incorporates the magnitude of the bandgap and its nature (direct or indirect), the shape of absorption spectra, the thickness of the absorber layer, the material-dependent non-radiative recombination losses, and the temperature. The standard solar spectrum (AM-1.5G), the absorption coefficient, thickness, and the electronic G0W0@PBE bandgap are thus used as inputs to evaluate the theoretical SLME of ASnX3 (A = Ca, Sr, Ba; X = S, Se) CPs at 293.15 K temperature.
In addition, the optical transition possibility from VBM to CBM for these CPs has been confirmed through the computation of transition dipole moment matrix (P) elements; its square (P2) gives the transition probability between the initial (VBM) and the final (CBM) state. Figure 4(a) shows the optically allowed dipole transition at -point for CaSnS3, and for other configurations, see Figure S7. Despite having an indirect electronic bandgap, BaSnS3 and BaSnSe3 exhibit an optically allowed dipole transition at their direct band edge. However, when is not the minimum bandgap of the materials (i.e., ), non-radiative recombination plays a vital role in the SLME calculation and the radiative recombinations vary with a factor , where , is the Boltzmann constant, and T is the temperature[44] (for details, see the Supplemental Material).
Next, the thickness dependence of SLME has been computed for all the ASnX3 (A = Ca, Sr, Ba; X = S, Se) compounds using BSE@G0W0@PBE method and plotted in Figure 4(b). It is clear that the SLME rises as thickness increases and eventually saturates beyond a certain thickness. The maximum SLME is found to be 31.20% for CaSnS3, which is consistent with the previously documented theoretical efficiency of 32.45% at 10 m thickness[14] and is higher than CH3NH3PbI3 (28.97% at 2 m)[28]. The highest SLME values for SrSnS3, BaSnS3, CaSnSe3, SrSnSe3, and BaSnSe3 compounds are calculated as 29.53%, 26.66%, 24.20%, 25.95%, and 26.09%, respectively. These values of SLME are favorable for the photovoltaic applications compared to transition metal-based traditional CPs and other halide perovskites[14, 78, 79, 80, 22, 15] (for details, see Table S14 of the Supplemental Material).
@@ III.7 Analysis of SCAPS-1D results:
To verify the photovoltaic potential of the investigated systems and our calculated SLME, we also conducted a detailed analysis using the Solar Cell Capacitance Simulator in 1 Dimension (SCAPS-1D). This simulation allowed us to evaluate key performance metrics and gain deeper insights into the efficiency and operational behavior of the devices. Here, the performance of PSC is evaluated by varying the thickness of the absorber layer while keeping the ETL, HTL, and FTO thicknesses constant. The absorber layer thickness is varied from 0.01 m to 1.2 m, and it found that the thickness of the perovskite layer has a significant impact on the performance of the PSC [see Figure 5(b)]. After reaching the optimum efficiency, the effect of the thickness of the perovskite layer seems to saturate, which agrees with our SLME calculations. The obtained results are summarized in the Table 7. Our findings indicate that the device simulation results are consistent with the SLME work, specifically for ASnS3 (A = Ca, Sr, Ba). On the other hand, the minor discrepancies between the device simulation results for ASnSe3 (A = Ca, Sr, Ba) and the SLME findings may be attributed to the materials used in the ETL, HTL, and FTO layers. In our study, SLME provides an idealized efficiency limit based solely on the material’s intrinsic optical properties, with its behavior improving for thicker films as absorption approaches a step function, aligning with the SQ limit. However, SCAPS-1D accounts for additional real-world effects, such as carrier transport, recombination losses, and device architecture, which can result in a performance decay for thicker films beyond 1 m due to increased recombination or reduced carrier collection efficiency. This distinction highlights the necessity of combining both methods to achieve a comprehensive assessment of the material’s optoelectronic potential, effectively bridging the gap between idealized theoretical predictions and practical device limitations. In general, SCAPS-1D proves to be a highly reliable software, as its simulation results confirm the high efficiency of the investigated materials, closely aligning with theoretical predictions.
@@
Overall, the study of all the above-discussed properties collectively shows that Se-based CaSnSe3 and SrSnSe3 demonstrate lower exciton binding energies and higher charge carrier mobilities than the rest of the investigated systems. However, their SLME is also relatively low. On the other hand, S-based CaSnS3 and SrSnS3 perovskites exhibit higher SLME despite having higher exciton binding energies and lower charge carrier mobilities compared to the other examined materials. This improved SLME is primarily attributed to the optimal bandgaps of CaSnS3 and SrSnS3 compared to their selenium (Se) counterparts, specifically from the perspective of solar cells. This reveals a trade-off between exciton binding energy, charge carrier mobility, and the bandgap of these materials. This suggests that integrating a combination of sulfur (S) and selenium (Se) atoms in these materials could be advantageous for potential application in solar cells, meriting further investigation.
@@ IV Conclusions:
In conclusion, we have carried out a comprehensive study to investigate the ground- and excited-state properties of post-transition metal Sn-based distorted chalcogenide perovskites (ASnX3; A = Ca, Sr, Ba, and X = S, Se) under the framework of state-of-the-art DFT combined with DFPT and MBPT (viz., GW and BSE). The mechanical properties confirm the stability, and Pugh’s and Poisson’s ratios reveal the ductile nature of these perovskites. The bandgaps calculated using G0W0@PBE method are in the range of 0.791.50 eV, and the small effective masses of electrons suggest good charge carrier mobility, which is advantageous for energy-harvesting properties. Furthermore, they have a high optical absorption coefficient ( 104 cm-1) and an optical bandgap extending from the near-infrared to the visible range, which is indicated by the BSE calculations. The exciton binding energies of these compounds, ranging from 0.04 to 0.23 eV, are similar to those of transition metal (Zr and Hf) based CPs. Also, these perovskites have smaller carrier-phonon coupling strengths than the conventional lead-based HPs as well as transition metal-based CPs, which leads to better polaron mobility for electrons (21.33416.02 cm2V-1s-1) and holes (7.02260.69 cm2V-1s-1). Additionally, the Fröhlich mesoscopic model implies that the charge-separated polaronic states have lower stability in comparison to bound excitons. Lastly, the SLME method forecasts that one could achieve the highest PCE of up to 31.20% by employing them, which is corroborated by conventional device (FTO/TiO2/ASnX3/Spiro-OMeTAD/Au) simulations using SCAPS-1D software. The enhanced properties of Sn-based CPs can be attributed to the distinct valence state of Sn in comparison to transition metal-based CPs. Overall, these results are anticipated to expedite the research and use of chalcogenide perovskites in optoelectronic applications, in general, and solar cell technology, in particular.
@@ Acknowledgements.
The authors would like to acknowledge the Council of Scientific and Industrial Research (CSIR), Government of India [Grant No. 3WS(007)/2023-24/EMR-II/ASPIRE] for financial support. S.A. would also like to acknowledge the CSIR, India [Grant No. 09/1128(11453)/2021-EMR-I] for the Senior Research Fellowship. S. D. thanks Prof. Marc Burgelman, Department of Electronics and Information Systems (ELIS), University of Gent, for providing the SCAPS-1D software. The authors acknowledge the High Performance Computing Cluster (HPCC) ’Magus’ at Shiv Nadar Institution of Eminence for providing computational resources that have contributed to the research results reported within this paper.References